A Complete Bibliography of Publications in Technology and Culture: 1950–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

30 September 2023 Version 1.00

Title word cross-reference

14th [Smi59]. 18th [Smi59]. 1955 [Bar59].

American [Bou59]. Appreciation [Dun59]. Atlantic [Hic59]. Automobile [Bou59].

B [Bou59]. Bern [Hic59]. Brief [Ano59b].

Cable [Hic59]. Centuries [Smi59]. Century [Con59]. Chain [Smi59]. Change [All59]. Commission [Bar59]. Culture [Bur59]. Current [All59].

Dibner [Hic59]. Dr. [Ano59a].

Excavations [Bar59]. Expression [Con59].

REFERENCES 2

Fielding [Ano59a, Dun59]. First [Bou59]. Forty [Bou59]. Hardware [Bur59]. Historical [Bar59]. History [Ano59b, Jon59]. Ideas [Jon59]. Improver [Mul59]. John [Bou59]. Khami [Bar59]. Mail [Smi59]. Making [Smi59]. Manufacturers [Bou59]. Memoriam [Ano59a]. Metallographic [Smi59]. Methods [Smi59]. Monuments [Bar59]. Natural [Bar59]. Nineteenth [Con59]. Note [Smi59]. Ogburn [Ano59a, Dun59]. Outlook [All59]. Preservation [Bar59]. R [Bar59]. Rae [Bou59]. Railway [Gei59]. Relics [Bar59]. Report [Bar59]. review [Bar59, Bou59, Hic59]. Rhodesia [Bar59]. Robinson [Bar59]. Ruins [Bar59]. Scientist [Mul59]. Skyscrapers [Con59]. Social [All59]. Society [Ano59b]. Southern [Bar59]. Start [Kra59]. Status [All59]. Sullivan [Con59]. Technology [All59, Con59, Jon59, Mul59, Ano59b]. Tools [Dru59]. Undertaken [Bar59]. William [Ano59a, Dun59]. Work [Dru59]. Years [Bou59].

References

Allen:1959:TSC

[All59] Francis R. Allen. Technology and social change: Current status and outlook. Technology and Culture, 1(1):48–59, Winter 1959. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

        https://muse.jhu.edu/pub/1/article/895237/pdf.

Anonymous:1959:MDW

[Ano59a] Anonymous. In memoriam: [Dr. William Fielding Ogburn]. Technology and Culture, 1(1):ii, Winter 1959. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/895246/pdf. See also [Dun59].

Anonymous:1959:SHT

[Ano59b] Anonymous. The Society for the History of Technology: a brief history. Technology and Culture, 1(1):106–108, Winter 1959. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895245/pdf.

Barnett:1959:KRR

[Bar59] H. G. Barnett. Khami Ruins. Report on Excavations Undertaken for the Commission for the Preservation of Natural and Historical Monuments and Relics, Southern Rhodesia. 1947–1955 by K. R. Robinson (review). Technology and Culture, 1(1):100–102, Winter 1959. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/895242/pdf.

Boulding:1959:AAM

[Bou59] K. E. Boulding. American Automobile Manufacturers: The First Forty Years by John B. Rae (review). Technology and Culture, 1 (1):104–105, Winter 1959. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895244/pdf.

Burlingame:1959:HC

[Bur59] Roger Burlingame. The hardware of culture. Technology and Culture, 1(1):11–19, Winter 1959. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/895233/pdf.

Condit:1959:SSE

[Con59] Carl W. Condit. Sullivan’s skyscrapers as the expression of Nineteenth Century technology. Technology and Culture, 1(1):78–93, Winter 1959. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895240/
pdf.

Drucker:1959:WT

[Dru59] Peter F. Drucker. Work and tools. Technology and Culture, 1(1):28– 37, Winter 1959. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
895235/pdf.

Duncan:1959:AWF

[Dun59] Otis Dudley Duncan. An appreciation of William Fielding Ogburn. Technology and Culture, 1(1):94–99, Winter 1959. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895241/pdf. See also [Ano59a].

Geise:1959:WR

[Gei59] John Geise. What is a railway? Technology and Culture, 1(1):68–77, Winter 1959. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895239/
pdf.

Hickernell:1959:ACB

[Hic59] L. F. Hickernell. The Atlantic Cable by Bern Dibner (review). Technology and Culture, 1(1):102–104, Winter 1959. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895243/pdf.

Jones:1959:IHT

[Jon59] Howard Mumford Jones. Ideas, history, technology. Technology and Culture, 1(1):20–27, Winter 1959. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/895234/pdf.

Kranzberg:1959:S

[Kra59] Melvin Kranzberg. At the start. Technology and Culture, 1(1):1–10, Winter 1959. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895232/
pdf.

Multhauf:1959:SIT

[Mul59] Robert P. Multhauf. The scientist and the “Improver” of technology. Technology and Culture, 1(1):38–47, Winter 1959. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/895236/pdf.

Smith:1959:MMC

[Smi59] Cyril Stanley Smith. Methods of making chain mail (14th to 18th centuries): a metallographic note. Technology and Culture, 1(1):60–67, Winter 1959. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/895238/
pdf.